# "I understand why I got this grade": Automatic Short Answer Grading with Feedback

### v1

@@ 4 Experiments
Building on the foundational work by Filighera et al. (2022) in their investigation, our primary objective is to conduct experiments to establish a baseline for our EngSAF dataset. We also seek to delve deeper into the impact of incorporating questions on the generated feedback and the assigned labels. Traditionally, in Automated Short Answer Grading (ASAG), assessments have focused solely on evaluating reference answers and student responses. However, Lv et al. (2021) challenge this convention by demonstrating that integrating questions into the evaluation process enhances the performance of traditional ASAG tasks. Therefore, our study aims to build upon these insights to explore the nuanced effects of incorporating questions in ASAG assessments, aiming to refine and improve the evaluation criteria and methodologies in this domain.
@@ 4.1 Experimental Setting
Fine-tuning Large Language Models (LLMs): We use the Llama-2 model Touvron et al. (2023), Mistral 7b model555https://mistral.ai/, which are fine-tuned to predict the output label for student responses, categorizing them as correct, partially correct, or incorrect and jointly providing feedback explaining the assigned output label. Furthermore, we have conducted this experiment using two distinct methodologies.
@@
The following models are used in this experiment.
llama-2-13B-chat
Mistral-7B-Instruct-v0.1 -
@@
Zero-shot experiment using chatGPT: For this experiment, we prompted ChatGPT to assign an output label along with feedback to each student’s response by evaluating its correctness compared to the reference answer for the ASAG task. Prompt details are present in the Techinical Appendix. We used ChatGPT API666https://openai.com/ to perform this experiment on EngSAF Dataset. The following model has been used in this experiment.
gpt-3.5-turbo-16k
Hyperparameters used for fine-tuning LLMs on EngSAF can be found in Technical Appendix 10.2. All experiments were performed on 2 NVIDIA A100-SXM GPU with 80 GiB of memory. Fine-tuning takes about 1 hour per epoch for training.
@@ 5 Results
Table 5 shows a majority baseline, Llama-2 fine-tuning, Mistral-7b fine-tuning, and ChatGPT zero-shot experiments results. The majority baseline contains the most occurring label and feedback from the EngSAF train set. The most common label is “correct response,” and the most common feedback is “Well done! You have answered the question correctly, covering all the required aspects.”. We conducted the ChatGPT experiment only on the unseen answers test set due to its zero-shot setting, as both unseen questions and answers are identical in the zero-shot setting.
@@
| ChatGPT | 41.1 | 24.0 | 7.4 | 31 | 11.3 | 28.8 | - | - | - | - | - | - |
@@ 5.1 Analysis
In this section, we delve into the insights drawn from the performance results of experiments on the EngSAF dataset for the ASAG task. The evaluation metrics used for labels are accuracy and macro-averaged F1. To evaluate the feedback, we measure the Rouge-2 Post (2018), SCAREBLEU 777https://pypi.org/project/sacrebleu/, METEOR Banerjee and Lavie (2005) and BERTSCORE Zhang et al. (2019) scores.
Table 5 shows that Mistral-7b significantly outperforms the majority baseline in both label and feedback metrics. However, a notable disparity exists in improvement between unseen questions and unseen answers. We observe a decrease of approximately 23% in accuracy for unseen questions compared to unseen answers. This underscores the challenge even fine-tuned models face when tasked with generalizing to new questions. Additionally, there is approximately 4% decrease in Bertscore for unseen questions compared to unseen answers, suggesting the necessity for new evaluation metrics that comprehensively assess text on the context level instead of the lexical level. Including questions in the input has led to a notable improvement in Bertscore, with a 10% increase for unseen answers and a 12% increase for unseen questions. Similarly, for label accuracy, we observe a 3.5% increase for unseen answers and a 7.3% increase for unseen questions in the test set. Technical Appendix includes qualitative analysis of test examples. Overall, the experiment suggests including questions for improvement in the predicted feedback by the model. The ChatGPT zero-shot experiment yields the lowest accuracy at 41.1%, even falling below that of the majority baseline, highlighting the complexity of the ASAG task and struggle of large language models like ChatGPT struggle in a zero-shot setting. Another contributing factor to this performance disparity could be the inherent bias present in teacher grading, which models can only learn when trained on such data. Interestingly, the smaller Mistral 7B model outperformed the LLaMA-2-13B model for ASAG, demonstrating the potential of more compact models in this task.
6 Deployment
@@
| Feedback Quality/Correctness | 4.5 |
| Emotional Impact | 4.9 |
For the real-world deployment, our fine-tuned ASAG model was integrated into an end-semester exam on ET 623 (Learning Analytics Course) at IIT Bombay for the 2024 academic year. The setup included the students enrolled in the course who consented to participate in the experiment. The end-sem exam includes 2 short-answer questions, each accompanied by the instructor’s correct/reference answer. We randomly sampled 25 student answers for each question and used our fine-tuned ASAG with Feedback model to predict the output label and feedback/ explanation for the predicted output. Each predicted output label was evaluated based on its correctness, where the subject matter expert checks whether the predicted output label is correct or not. Each predicted feedback is analyzed in terms of Feedback Correctness/ Accuracy and Emotional Impact as discussed in the Quality Estimate section 3.5. Each aspect is scored on a scale (1-5), with a high score indicating a better response.
@@
Table 6 presents the average scores across all the designed aspects used to measure the reliability of the proposed dataset. Upon examining the table, it is evident that the subject matter evaluation scores greater than 4.5 for both Feedback Quality/Correctness and the Emotional Impact aspect. This demonstrates the reliability and effectiveness of the ASAG model in real-world scenarios. Additionally, the model’s predicted output label achieved an accuracy of 92.5%, further showcasing its performance and reliability. Technical Appendix contains more details about deployment.
7 Summary, Conclusion, and Future Work
This study presents a novel and extensive dataset for grading short answers across multiple engineering domains. Our dataset addresses the need for a standardized evaluation platform by covering diverse subjects and a wide range of short answers. We meticulously curated and annotated this dataset to ensure its quality and applicability. Using the Label Aware Synthetic Feedback Generation (LASFG) strategy we have synthetically incorporated feedback in the EngSAF dataset. Through experimentation, we benchmarked the dataset’s performance using different LLM’s fine-tuning and chatGPT zero shot experiments.
In conclusion, our research endeavors culminated in successfully creating and evaluating a cutting-edge multi-domain short answer grading dataset. The dataset’s diverse content and meticulous annotations provide a solid platform for training and assessing grading models across various subject domains. Our benchmarking experiments showcased promising results, indicating that our dataset has the potential to significantly improve the quality and reliability of ASAG systems. Additionally, we deployed the ASAG system in a real-world setting for an IITB course exam, demonstrating its practical applicability. This work marks a crucial step forward in educational technology, equipping educators and researchers with a valuable resource to boost advancements in automated assessment techniques.
While this research presents a novel dataset for effective short answer grading along with feedback, there are several possibilities for future investigation. One area of focus could be the refinement and expansion of our dataset to include more nuanced and complex short answers, enabling us to accommodate a broader range of grading scenarios. Baseline figures could be improved by expansion or augmentation of the dataset. External knowledge inclusion through knowledge graphs could be explored in this direction.
8 Limitations
Our study, titled “I understand why I got this grade": Automatic Short Answer Grading with Feedback" has provided valuable insights into the multi-domain ASAG. However, it’s important to acknowledge certain limitations. The synthetic nature of the generated feedback may introduce biases or might contain inconsistencies, factual inaccuracies, or overly generic content, that could impact the implications of the results. One key finding from table 3 is an imbalance in output labels. Specifically, Output label Zero (0) is less frequent compared to labels One (1) and Two (2). This imbalance could impact the performance of automatic grading models trained on this dataset. Despite our efforts to curate diverse short-answer responses across multiple engineering domains, the dataset’s overall size might limit the complexity and depth of models that can be trained on it. Variations in data distribution, language patterns, and task requirements across different domains and contexts may impact the application of our results to real-world scenarios. While our dataset’s strength lies in its coverage of multiple engineering domains, it’s important to note that the range of domains might not be exhaustive. Researchers focusing on specific domains or aiming to enhance domain-specific performance may need to consider domain-specific or subject-specific fine-tuning or gather additional data.
9 Ethical Statement
The EngSAF dataset contains questions, reference answers, and student answers along with the output label assigned for each student answer. This dataset is taken from actual exams and quizzes conducted at the university. We are committed to protecting those student’s privacy and anonymity. Hence, no student-specific information is disclosed within the dataset. The process of collecting and curating this dataset adhered to the ethical norms of informed consent. Prior to any contribution, instructors of each course were provided details concerning the data collection’s intent, the utilization of their questions and responses, and any potential implications. While the ASAG models created with this dataset offer great potential in educational assessment, they must be used properly and ethically. As with any AI technology, its impact on education must be regularly monitored and evaluated. Ensuring an appropriate balance between the advantages derived from automation and safeguarding the educational experience and human judgement is of utmost importance.

### v2

@@ 4 Experiments
Following Filighera et at.’s [12] methodology, our primary objective is to establish a baseline for our EngSAF dataset. We also assess the impact of incorporating questions on the generated feedback and the assigned labels, by performing experiments in two different settings, namely with_question and without_question. Traditionally, in Automatic Short Answer Grading (ASAG), assessments have focused solely on evaluating reference answers and student responses. However, Lv et al. [19] challenge this convention by demonstrating that integrating questions into the evaluation process improves the performance of traditional ASAG tasks.
@@ 4.1 Experimental Setting
Fine-tuning Large Language Models (LLMs): We use the Llama-2 model [35], Llama-3.1 model [9], Mistral 7b model666https://mistral.ai/, which are fine-tuned to predict the output label for student responses, categorizing them as correct, partially correct, or incorrect and jointly providing feedback explaining the assigned output label. Furthermore, we have conducted this experiment using two distinct methodologies.
@@
The models used in this experiment are llama-3.1-8B-Instruct-chat, llama-2-13B and Mistral-7B-Instruct-v0.1.
@@
Zero-shot experimentation: For this experiment, we prompt ChatGPT777https://openai.com/ and DeepSeek [14] to assign an output label along with feedback to each student’s response by evaluating its correctness compared to the reference answer for the ASAG task. The following models are used: gpt-4o and DeepSeek-R1
All experiments are performed on 2 NVIDIA A100-SXM GPU with 80 GB of memory. Fine-tuning takes around 2 hours per epoch for training.
@@ 5 Results
| Zero-shot experiments | ||||||||||||
| ChatGPT-4o | 48.9 | 46.9 | 3.64 | 22.3 | 5.2 | 20.8 | - | - | - | - | - | - |
| DeepSeek-R1 | 47.4 | 46.7 | 4.1 | 23.6 | 6.5 | 18.0 | - | - | - | - | - | - |
Table 5 shows a majority baseline, fine-tuning experiments, and zero-shot experiments results. The majority baseline contains the most occurring label and feedback from the EngSAF train set. The most common label is “correct response,” and the most common feedback is “Well done! You have answered the question correctly, covering all the required aspects.”.
@@ 5.1 Analysis
In this section, we delve into the insights drawn from the performance results of experiments on the EngSAF dataset for the ASAG task. Following Filighera et at.’s [12] evaluation methodology, we measure the accuracy and macro-averaged F1 for output labels and the ROUGE-2 [26], SACREBLEU888https://pypi.org/project/sacrebleu/, METEOR [6] and BERTScore [37] for the feedback part. For all metrics, a higher score indicates better results.
As per Table 5, Mistral-7b significantly outperforms the majority baseline in both output label and feedback metrics. However, we observe a significant performance gap between unseen answers and unseen questions, with accuracy dropping by approximately 23% for unseen questions and by 4% in BERTScore, suggesting the necessity for new evaluation metrics that comprehensively assess text on the context level instead of the lexical level. The above two observations highlight the challenge that even fine-tuned models face when generalizing to new questions or domains (UQ). In the case of unseen answers (UA), other responses to the same question are present in the training data, making generalization easier compared to entirely new questions. A similar pattern can be observed for other fine-tuned models. Notably, including questions in the input significantly improves accuracy, F1 score, and BERTScore. For example, in our experiments with Llama-2 and Mistral-7B, incorporating the question led to a 10% improvement in BERTScore for unseen answers and a 12% improvement for unseen questions.
Our zero-shot experiments with ChatGPT and DeepSeek-R1 yield an accuracy of 48.9% and 47.4%, respectively, falling below the fine-tuning baselines, highlighting the complexity of the ASAG task. We observe a similar performance gap in the feedback evaluation part, suggesting that fine-tuning plays a crucial role in capturing complex grading patterns. Another contributing factor to this performance disparity of zero-shot experiments in comparison to fine-tuned experiments could be the inherent bias present in teacher grading, which models can only learn when trained on such data.
6 Real-World Deployment
For real-world deployment, we integrated our best-performing fine-tuned ASAG model for automatic evaluation of an end-semester exam of a course at a reputed institute. The end-semester examination included two short-answer questions, each accompanied by the instructor’s correct/reference answer. We randomly sampled 25 student answers for each question and used our fine-tuned model to predict the output label and feedback/ explanation for the predicted output. To ascertain the quality of the generated output, we employed three Subject Matter Experts (SMEs) to evaluate the predicted output and feedback. The SMEs were PhD students who were teaching assistants (TAs) of that course. Each predicted output label was evaluated based on its correctness, where the subject matter expert checks whether the predicted output label is correct or not. Each predicted feedback is analyzed in terms of Feedback Quality/Correctness (FC) and Emotional Impact (EI) as discussed in the Quality Estimate section 3.4. Each aspect is scored on a scale (1-5), with a high score indicating a better response.
@@ 6 Real-World Deployment
| Feedback Quality/Correctness (FC) | 4.5 |
| Emotional Impact (EI) | 4.9 |
@@
Table 6 presents the average scores across all the aspects used to measure the reliability of our ASAG model. Upon examining the table, it is evident that the subject matter evaluation scores are greater than 4.5 for both FC and the EI aspect. This demonstrates the reliability and effectiveness of the ASAG model in real-world scenarios. Additionally, the model’s predicted output label achieved an accuracy of 92.5%, further showcasing its performance and reliability. By providing accurate labels and feedback at scale, our approach helps reduce the manual workload for teachers and instructors, making the grading process more efficient and scalable.
7 Conclusion
This work introduces a novel and extensive dataset for automatic short answer grading (ASAG) across multiple engineering domains. Addressing the need for a standardized evaluation platform, our dataset encompasses diverse subjects and a wide range of short-answer responses. We meticulously curated and annotated the dataset to ensure its quality, consistency, and applicability to real-world grading scenarios. We propose Label-Aware Synthetic Feedback Generation (LASFG)—a method that enables the transformation of any traditional ASAG dataset into one that incorporates meaningful synthetic feedback. Additionally, we benchmark our dataset using fine-tuned large language models (LLMs) and conduct zero-shot evaluations, providing a comprehensive assessment of its effectiveness.
Furthermore, we deployed our ASAG system in a real-world course exam at a prestigious institution, showcasing its practical utility in educational settings. Human evaluation showcases the effectiveness of the ASAG system by achieving scores more than 4.5 out of 5 on Feedback Quality/Correctness (FC) and Emotional Impact (EI). Additionally, the model demonstrates its ability to correctly detect an answer with an accuracy of 92.5%. This work represents a significant milestone in educational technology, offering educators and researchers a valuable tool to drive innovation in automated assessment methods.
Future extensions of this work could focus on expanding the dataset to include more complex short-answer responses, enabling broader grading scenarios. Additionally, incorporating the evaluation of diagrams and images within student responses could further enhance the system’s capabilities, paving the way for more comprehensive automated assessment solutions.